Please I would appreciate anyone's help. I am stuck on this problem.
Solve by elimination or substitution. For all three variables a, b and c. Use all three equations.
5a-5b+5c=25
-4a+4b-2c=-12
-a+ b-5c=-21

Answer by TimothyLamb(4379)

5a-5b+5c=25
-4a+4b-2c=-12
-a+ b-5c=-21
---
put the system of linear equations into standard form
---
substitute:
x for a
y for b
z for c
---
5x-5y+5z=25
-4x+4y-2z=-12
-x+ y-5z=-21
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
This linear system is singular. It has no solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php